# When Your Boss Is an AI Bot: Exploring Opportunities and Risks of Manager Clone Agents in the Future Workplace

### 24 Sep 2025 v2

@@ Abstract.
As Generative AI (GenAI) becomes increasingly embedded in the workplace, managers are beginning to create Manager Clone Agents—AI-powered digital surrogates that are trained on their work communications and decision patterns to perform managerial tasks on their behalf. To investigate this emerging phenomenon, we conducted six design fiction workshops (n = 23) with managers and workers, in which participants co-created speculative scenarios and discussed how Manager Clone Agents might transform collaborative work. We identified four potential roles that participants envisioned for Manager Clone Agents: proxy presence, informational conveyor belt, productivity engine, and leadership amplifier, while highlighting concerns spanning individual, interpersonal, and organizational levels. We provide design recommendations envisioned by both parties for integrating Manager Clone Agents responsibly into the future workplace, emphasizing the need to prioritize workers’ perspectives, strengthen interpersonal bonds, and enable flexible clone configuration.
Manager Clone Agents, AI Agent, Future of Work, Generative AI, Workplace, Human–AI Collaboration
conference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYccs: Human-centered computing Computer supported cooperative workccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Recent advances in AI have given rise to the phenomenon of managers “cloning themselves” to delegate routine and strategic work such as attending meetings, exchanging information with employees and stakeholders, and assisting in decisions (Forsdick, 2025; Raconteur, 2023; Gani, 2025). Unlike algorithmic systems that primarily optimize decisions or workflows (Lee et al., 2015), Manager Clone Agents replicate a manager’s symbolic, social, and communicative presence. They act as digital clones (Liang, 2025; Lee et al., 2023), mimicking a manager’s appearance, behaviors, and communication style, and can also evolve into functional delegates (Cheng et al., 2025; The Journey Platform, 2025) that perceive context, reason through tasks, and execute complex actions without intervention.
@@
Yet alongside this momentum lie pressing concerns. Because managers shape decision-making, resource distribution, and workplace culture (Mintzberg et al., 1973), delegating aspects of their authority, whether symbolically or functionally, to AI agents raises difficult questions of trust, fairness, and accountability. Workers worry about displacement and the erosion of authentic communication (Bonos and Abril, 2025). These tensions make it urgent to study how such agents alter leadership, collaboration, and organizational dynamics, and to identify pathways for their responsible design and adoption.
Existing HCI studies have begun to examine how AI Clones, personalized AI mimics that approximate individuals given limited training data (Liang, 2025), coexist with their human originals, often focusing on questions of identity, self-perception and other interpersonal relationships (Hwang et al., 2024; Huang et al., 2025; Wang et al., 2025; Lee et al., 2023). Building on this work, we use the term Manager Clone Agents to refer broadly to a family of emerging systems, ranging from symbolic clones that replicate presence to functional delegates that automate managerial tasks. Positioned at the intersection of authority and collaboration, they combine task-oriented delegation with the symbolic and communicative authority of human managers. This dual capacity makes them distinct from prior managerial technologies (Lee et al., 2015; Wolf and Stock-Homburg, 2025; Sowa et al., 2020), but we know little about how their human-like qualities will reshape teamwork within organizations.
@@
To explore these questions, we conducted a series of design fiction workshops (Lyckvi et al., 2018; Muller et al., 2020) with managers and workers, asking participants to imagine organizational futures where Manager Clone Agents were widely adopted. Design fiction, a method widely used in HCI, was chosen because Manager Clone Agents remain speculative technologies in this moment for most managers and workers: it allows participants to engage with fictional yet plausible scenarios to articulate hopes, conflicts, and concerns without being constrained by today’s technical limits. In this paper, we define “managers” as organizational leaders at multiple levels, including executives, directors, and middle managers, and “workers” as their employees and direct reports who experience the effects of managerial authority. In the workshops, both groups created scenarios of Manager Clone Agents use and then co-reflected on the opportunities, tensions, and uncertainties these agents might generate in organizational life.
Our findings revealed a dual character to Manager Clone Agents. On one hand, participants envisioned supportive roles, acting as proxy presences in meetings, conveying information across organizational layers, automating routine tasks, and amplifying managerial guidance (RQ1). On the other hand, these same roles raised risks across multiple levels: at the individual level, from different positions both managers and workers experienced shared anxieties about accountability and irreplaceability; at the interpersonal level, both sides pointed to fragile trust, blurred authenticity, and reduced informal contact; and at the organizational level, participants worried that efficiency gains could flatten hierarchies and weaken belonging (RQ2). Participants therefore stressed that responsible design must foreground workers’ voices, preserve communication and social cohesion, and allow trust to develop gradually through clear boundaries on what Manager Clone Agents can and cannot do (RQ3).
This study makes three key contributions to HCI and CSCW community:
  • First, we present empirical vision, shaped together by managers and workers of potential roles that emerging Manager Clone Agents could undertake and surface the opportunities and risks these roles create across individual, interpersonal, and organizational levels in future workplaces.
@@
  • Finally, we surface design directions for developing Manager Clone Agents that build trust and evolve collaborations within diverse team and organizational contexts.
@@ 2. Background: Emerging Manager Clone Agents in Workplaces
Recent advances in artificial intelligence have enabled the rise of Manager Clone Agents. Executives are beginning to use these tools to manage heavy schedules, from financial briefings to routine worker inquiries. As early as 2021, Nvidia CEO Jensen Huang released an AI avatar powered by Omniverse Avatar (NVIDIA, 2022). Since then, startups such as Personal AI [^1] https://www.personal.ai/ , Delphi [^2] https://www.delphi.ai/ , and Tavus [^3] https://www.tavus.io/ have built platforms that train avatars on leaders’ speeches and meetings to create chatbots, voice agents, and video avatars. Investment has followed quickly: Delphi raised \$16M in June 2025 (Ladjevardian, 2025), HeyGen \$60M (Team, 2025), and ElevenLabs \$180M at a \$3.3B valuation (Mehta and Lunden, 2025), while Synthesia reached \$2.1B (Lunden, 2025). High-profile leaders such as Linkedin co-founder Reid Hoffman, Zoom CEO Eric Yuan, and Khosla Ventures managing director Keith Rabois have also begun experimenting with these systems (HeyGen, 2024; Patel, 2024; Morrone, 2025), using them for financial updates (Bort, 2025; Peters, 2025), meeting attendance (Boyle, 2025), or routine staff communication (Personal AI, 2025; Ladjevardian et al., 2025; Langley, 2025; Cuthbertson, 2023). Adoption of AI agents in workplaces is also accelerating. A PwC survey of 308 U.S. managers in April 2025 reported that 79% of companies already use AI agents, and 88% plan to expand budgets within a year (PwC, 2025). Falling costs are lowering entry barriers (Fifer and Feagans, 2025), positioning Manager Clone Agents to spread from executives to middle and line managers in the near future (DigitalDefynd, 2025). Yet concerns remain. Because managerial roles shape decisions, resources, and culture, delegating them to AI raises questions of trust, accountability, and fairness (Sowa et al., 2020). Scholars warn that these systems may reinforce executive detachment (Morrone, 2025), while workers fear job loss and erosion of authentic communication (Writer, 2025). These tensions highlight the need to better understand how Manager Clone Agents affect collaboration and organizational life.
@@ 3. Related Work
First, we review scholarship on managerial roles in the workplace, which highlights the complex mix of decision-making, coordination, and relational care that define managerial labor (see subsection 3.1). Second, we discuss HCI and CSCW work that examines the role of technologies in supporting managerial work, from algorithmic management to collaborative robots and AIs (see subsection 3.2). Third, we examine early work on Clone Agents studies, focusing on their potential to extend human presence while raising new concerns around authenticity (see subsection 3.3).
@@ 3.1. Managerial Roles in the Workplace
Managers occupy a distinctive position in organizational life. Unlike most employees, they are responsible for abstract decision-making, conflict coordination, and emotional motivation, which make their tasks qualitatively different from those of their subordinates (Dler and Tawfeq, 2021). Classic organizational theory frames managerial work as a constellation of intertwined roles: interpersonal (figurehead, leader, liaison), informational (monitor, disseminator, spokesperson), and decisional (entrepreneur, disturbance handler, resource allocator, negotiator) (Mintzberg et al., 1973). Yet these responsibilities accumulate into heavy burdens. Managers already face fragmented routines of meetings, emails, and constant communication that limit time for reflection and strategy (Stray and Moe, 2020; Dler and Tawfeq, 2021). Recent shifts in digital technologies that multiply channels and complicate coordination (Andersone et al., 2022), along with the demands of remote and hybrid work (Morrison-Smith and Ruiz, 2020; Choudhury et al., 2023) have further amplified these strains.
Beyond the structural demands of the role, effective management depends on the individual ways managers enact their responsibilities. Leadership research highlights how different decision-making styles (e.g., autocratic, democratic) produce divergent effects on team performance, satisfaction, and trust (Yukl and Fleet, 1992). Some managers also construct personal narratives that align organizational goals, thereby legitimizing their authority (Rostron, 2022). These perspectives underscore that effectiveness in management hinges less on a fixed formula and more on how managers’ personal identities and the relational dynamics they cultivate. Organizational efforts to ease workload, such as appointing chiefs of staff, executive assistants, or project management offices, reinforce this point: the success of such delegation hinges on trust, specifically on whether the decision-making comes from trusted internal leaders or less familiar external consultants (Meagher and Wait, 2020; Tamkin, 2014).
@@ 3.3. Clone Agents in Workplace
With the rise of GenAI, AI agents have emerged as a distinct category of intelligent digital assistants, capable of autonomously performing tasks on behalf of users and mimicking aspects of human interaction (Park et al., 2023, 2024; Li et al., 2024). A growing stream of research explores their role in teamwork, where AI agents can support managerial functions ranging from scheduling (Cranshaw et al., 2017) to decision-making (Li et al., 2025). Systems such as Cobot-style workflow assistants (Sowa et al., 2020) illustrate how agents can reduce leaders’ follow-up burdens. Other studies highlight how conversational agents can facilitate group discussions, such as encourage equal participation (Do et al., 2022; Kim et al., 2020), and team cohesion (Benke et al., 2020). These developments signal the growing managerial relevance of AI agents in collaborative contexts.
Building on this trajectory, Clone Agents represent a special subtype of AI agents. Unlike generic digital assistants, Clone Agents are designed to replicate not only an individual’s decision logic (Truby and Brown, 2021) but also their voice, appearance, and conversational style (Dobre et al., 2022; Leong et al., 2024; Shamekhi et al., 2018). Scholars have begun to conceptualize these entities as AI clones—digital representations trained on personal data that can interact with others as if they were the real-world individual (Lee et al., 2023). Early HCI and CSCW systems hinted at this direction: Time-Travel Proxy enabled asynchronous participation through pre-recorded video (Tang et al., 2012), and Lee & Takayama showed how mobile remote presence let remote workers “live and work” with colleagues almost as if physically there (Lee and Takayama, 2011). More recent work extends these ideas: Dittos created embodied proxies for real-time meetings (Leong et al., 2024), and Cheng et al. studied how delegating conversational autonomy to agents affects multitasking in group interactions (Cheng et al., 2025).
@@ 4.1. Participants
ID Sex Age Work Experience Industry Team
W1 Female 28 1–2 years Design Product Team
M1 Male 25 1–2 years Renewable Energy Risk Management Team
W2 Female 24 1–2 years Automotive Manufacturing Design Team
M2 Female 29 3–5 years Automotive Manufacturing Design Team
M3 Male 25 3–5 years Film and Television Production Technical Production Team
M4 Male 24 1–2 years Media Event Planning Team
W3 Female 24 0–1 year Internet Data Production Team
W4 Female 26 0–1 year Technology Research Team
M5 Female 30 1–2 years Education Product Team
M6 Male 28 3–5 years Internet Cross-Functional Team
W5 Male 25 0–1 year Technology and Education Research Team
W6 Female 26 0–1 year Internet Product Team
M7 Female 27 1–2 years Government Agencies Project Execution Team
W7 Male 28 3–5 years Digital Media Editorial Team
W8 Male 27 3–5 years Healthcare Design Team
M8 Female 30 3–5 years Professional Development Cross-functional Team
M9 Female 31 5–10 years Marketing Marketing Team
M10 Male 32 5–10 years Foundation Model Infrastructure Software QA Team
M11 Male 45 ¿10 years Information Technology Product Team
W9 Female 25 1–2 years Racing Equipment Product Team
M12 Male 39 5–10 years Design Communication Design Team
M13 Female 33 5–10 years Technology UX/UI Design Team
W10 Female 27 3–5 years Design UX/UI Design Team
@@
Our sample included 13 managers (M1 to M13) and 10 workers (W1 to W10). Participants ranged in age from 24 to 45 years across sectors. Notably, some workers in our sample had relatively long professional experience, while some managers were comparatively early in their careers. This reflected the fact that managerial status was not strictly tied to years of employment but to current responsibilities: in some organizations, employees became team leads or middle managers within just a few years, while in others, individuals with longer tenure continued to work as direct contributors under managerial supervision. Role categorization in our study therefore followed participants’ self-identification based on their formal position and everyday duties, rather than professional tenure alone. During the workshops, participants were asked to fully inhabit the roles they currently held within their company or team, drawing on their lived experiences as managers or workers to co-create and reflect on speculative scenarios of imagined futures about Manager Clone Agents.

### 29 Jan 2026 v3

@@ Abstract.
As Generative AI (GenAI) becomes increasingly embedded in the workplace, managers are beginning to create Manager Clone Agents—AI-powered digital surrogates trained on their work communications and decision patterns to perform managerial tasks on their behalf. To investigate this emerging phenomenon, we conducted six design fiction workshops (n = 23) with managers and workers, in which participants co-created speculative scenarios and discussed how Manager Clone Agents might transform collaborative work. We identified four potential roles that participants envisioned for Manager Clone Agents: proxy presence, informational conveyor, productivity engine, and leadership amplifier, while highlighting concerns spanning individual, interpersonal, and organizational levels. We provide design recommendations envisioned by both parties for integrating Manager Clone Agents responsibly into the future workplace, emphasizing the need to prioritize workers’ perspectives and nurture interpersonal bonds while also anticipating alternative futures that may disrupt managerial hierarchies.
Manager Clone Agents, AI Agent, Future of Work, Generative AI, Future Workplace, Human–AI Collaboration
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790987isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Computer supported cooperative workccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Recent advances in AI have given rise to the phenomenon of managers “cloning themselves” to delegate routine and strategic work such as attending meetings, exchanging information with employees and stakeholders, and assisting in decisions (Forsdick, 2025; Raconteur, 2023; Gani, 2025). Unlike algorithmic systems that primarily optimize decisions or workflows (Lee et al., 2015), Manager Clone Agents replicate a manager’s symbolic, social, and communicative presence. They act as digital clones (Liang, 2025; Lee et al., 2023), mimicking a manager’s appearance, behaviors, and communication style, and can also serve into functional delegates (Cheng et al., 2025; The Journey Platform, 2025) that perceive context, reason through tasks, and execute complex actions without intervention (see Figure 1).
@@
Yet, alongside this momentum lie pressing concerns. Because managers shape decision-making, resource distribution, and workplace culture (Mintzberg, 1973), delegating aspects of their authority, whether symbolically or functionally, to AI agents raises difficult questions of trust, fairness, and accountability. Workers worry about displacement and the erosion of authentic communication (Bonos and Abril, 2025). These tensions make it urgent to study how such agents alter leadership, collaboration, and organizational dynamics, and to identify pathways for their responsible design and adoption.
Existing HCI studies have begun to examine how AI Clones (personalized AI mimics that approximate individuals given limited training data (Liang, 2025)) coexist with their human originals, often focusing on questions of identity, self-perception and other interpersonal relationships (Hwang et al., 2024; Huang et al., 2025; Wang et al., 2025; Lee et al., 2023). Building on this work, we use the term Manager Clone Agents to refer broadly to a family of emerging systems, ranging from symbolic clones that replicate presence to functional delegates that automate managerial tasks. Positioned at the intersection of authority and collaboration, they combine task-oriented delegation with the symbolic and communicative authority of human managers. This dual capacity makes them distinct from prior managerial technologies (Lee et al., 2015; Wolf and Stock-Homburg, 2025; Sowa et al., 2020), but we know little about how their human-like qualities will reshape teamwork within organizations.
@@
To explore these questions, we conducted a series of design fiction workshops (Lyckvi et al., 2018; Muller et al., 2020; Ma et al., 2025) with managers and workers, asking participants to imagine organizational futures where Manager Clone Agents were widely adopted. Design fiction, a method widely used in HCI, was chosen because Manager Clone Agents remain speculative technologies at this moment for most managers and workers: it allows participants to engage with fictional scenarios to articulate hopes, conflicts, and concerns without being constrained by today’s technical limits. In this paper, we define “managers” as organizational leaders at multiple levels, including executives, directors, and middle managers, and “workers” as their employees or direct reports who experience the effects of managerial authority. In the workshops, both groups created scenarios of Manager Clone Agents use and then co-reflected on the opportunities, tensions, and uncertainties these agents might generate in organizational life.
Our findings reveal a fundamental tension between the promise and risks of Manager Clone Agents. On the one hand, participants envisioned supportive roles, acting as proxy presences in meetings, conveying information across organizational layers, automating routine tasks, and amplifying managerial guidance (RQ1). On the other hand, these same roles raised risks across multiple levels: at the individual level, from different positions both managers and workers experienced shared anxieties about accountability and irreplaceability; at the interpersonal level, both sides pointed to fragile trust, blurred authenticity, and reduced informal contact; and at the organizational level, participants worried that efficiency gains could flatten hierarchies and weaken belonging (RQ2). Participants, therefore, stressed that responsible design must foreground workers’ voices, preserve communication and social cohesion, and allow trust to develop gradually through clear boundaries on the capabilities of manager clone agents, while remaining attentive to speculative alternative futures that can affect managerial authority (RQ3).
This study makes three key contributions to the HCI and CSCW community:
  • First, we present an empirical vision of potential roles that emerging Manager Clone Agents could undertake. Drawing on perspectives from both managers and workers, we surface the opportunities and risks these roles create across individual, interpersonal, and organizational levels in future workplaces.
@@
  • Finally, we surface design directions for developing Manager Clone Agents that build trust and support collaborations within diverse team and organizational contexts, including those challenging the current hierarchies.
@@ 2. Background: Emerging Manager Clone Agents in Workplaces
Recent advances in artificial intelligence have enabled the rise of Manager Clone Agents. Executives are beginning to use these tools to manage heavy schedules, from financial briefings to routine worker inquiries. As early as 2021, Nvidia CEO Jensen Huang released an AI avatar powered by Omniverse Avatar (NVIDIA, 2022). Since then, startups such as Personal AI [^1] https://www.personal.ai/ , Delphi [^2] https://www.delphi.ai/ , and Tavus [^3] https://www.tavus.io/ have built platforms that train avatars on leaders’ speeches and meetings to create chatbots, voice agents, and video avatars. Investment has followed quickly: Delphi raised \$16M in June 2025 (Ladjevardian, 2025), HeyGen \$60M (Team, 2025), and ElevenLabs \$180M at a \$3.3B valuation (Mehta and Lunden, 2025), while Synthesia reached \$2.1B (Lunden, 2025). High-profile leaders such as Linkedin co-founder Reid Hoffman, Zoom CEO Eric Yuan, and Khosla Ventures managing director Keith Rabois have also begun experimenting with these systems (HeyGen, 2024; Patel, 2024; Morrone, 2025), using them for financial updates (Bort, 2025; Peters, 2025), meeting attendance (Boyle, 2025), or routine staff communication (Personal AI, 2025; Ladjevardian et al., 2025; Langley, 2025; Cuthbertson, 2023). The adoption of AI agents in workplaces is also accelerating. A PwC survey of 308 U.S. managers in April 2025 reported that 79% of companies already use AI agents, and 88% plan to expand budgets within a year (PwC, 2025). Falling costs are lowering entry barriers (Fifer and Feagans, 2025), positioning Manager Clone Agents to spread from executives to middle and line managers in the near future (DigitalDefynd, 2025). Yet concerns remain. Because managerial roles shape decisions, resources, and culture, delegating them to AI raises questions of trust, accountability, and fairness (Sowa et al., 2020). Employees expressed concerns about the erosion of authentic communication, job loss, and the widening gap between executive privilege and labor precarity (Writer, 2025; Morrone, 2025). These tensions highlight the need to better understand how Manager Clone Agents affect the future workplace.
@@ 3. Related Work
First, we review the scholarship on managerial roles in the workplace, which highlights the complex mix of decision-making, coordination, and relational care that define managerial labor (see subsection 3.1). Second, we discuss HCI and CSCW work that examines the role of technologies in supporting managerial work, from algorithmic management to collaborative robots and AIs (see subsection 3.2). Third, we examine the early work on Clone Agents, focusing on their potential to extend human presence while raising new concerns about identity and authenticity (see subsection 3.3).
@@ 3.1. Managerial Roles in the Workplace
Managers occupy a distinctive position in organizational life. Unlike most employees, they are responsible for abstract decision-making, conflict coordination, and emotional motivation, which makes their tasks qualitatively different from those of their subordinates (Dler and Tawfeq, 2021). Classic organizational theory frames managerial work as a constellation of intertwined roles: interpersonal (figurehead, leader, liaison), informational (monitor, disseminator, spokesperson) and decisional (entrepreneur, disturbance handler, resource allocator, negotiator) (Mintzberg, 1973). Yet these responsibilities accumulate into heavy burdens. Managers already face fragmented meetings, emails, and constant communication that limit time for reflection and strategy (Stray and Moe, 2020; Dler and Tawfeq, 2021). Recent shifts in digital technologies that multiply channels and complicate coordination (Andersone et al., 2022), along with the demands of remote and hybrid work (Morrison-Smith and Ruiz, 2020; Choudhury et al., 2023) have further amplified these strains.
Beyond the structural demands of the role, effective management depends on the individual ways managers enact their responsibilities. Leadership research highlights how different decision-making styles (e.g., autocratic, democratic) produce divergent effects on team performance, satisfaction, and trust (Yukl and Fleet, 1992). Some managers also construct personal narratives that align organizational goals, thereby legitimizing their authority (Rostron, 2022). These perspectives underscore that effectiveness in management hinges less on a fixed formula and more on how managers’ personal identities and the relational dynamics they cultivate. Organizational efforts to ease workload, such as appointing chiefs of staff, executive assistants, or project management offices, reinforce this point: the success of such delegation hinges on trust, specifically on whether decision-making comes from trusted internal leaders or less familiar external consultants (Meagher and Wait, 2020; Tamkin, 2014).
@@ 3.3. Clone Agents in Workplace
With the rise of GenAI, AI agents have emerged as a distinct category of intelligent digital assistants, capable of autonomously performing tasks on behalf of users and mimicking aspects of human interaction (Park et al., 2023, 2024; Li et al., 2024). A growing stream of research explores their role in teamwork, where AI agents can support managerial functions ranging from scheduling (Cranshaw et al., 2017) to decision-making (Li et al., 2025). Systems such as Cobot-style workflow assistants (Sowa et al., 2020) illustrate how agents can reduce leaders’ follow-up burdens. Other studies highlight how conversational agents can facilitate group discussions, such as encourageing equal participation (Do et al., 2022; Kim et al., 2020), and team cohesion (Benke et al., 2020). These developments signal the growing managerial relevance of AI agents in collaborative contexts.
Building on this trajectory, Clone Agents represent a special subtype of AI agents. Unlike generic digital assistants, Clone Agents are designed to replicate not only an individual’s decision logic (Truby and Brown, 2021) but also their voice, appearance and conversational style (Dobre et al., 2022; Leong et al., 2024; Shamekhi et al., 2018). Scholars have begun to conceptualize these entities as AI clones—digital representations trained on personal data that can interact with others as if they were a real-world individual (Lee et al., 2023). Early HCI and CSCW systems hinted at this direction: Time-Travel Proxy enabled asynchronous participation through pre-recorded video (Tang et al., 2012), and Lee & Takayama showed how mobile remote presence let remote workers “live and work” with colleagues almost as if physically there (Lee and Takayama, 2011). More recent work extends these ideas: Dittos created embodied proxies for real-time meetings (Leong et al., 2024), and Cheng et al. studied how delegating conversational autonomy to agents affects multitasking in group interactions (Cheng et al., 2025).
@@ 4.1. Participants
ID Role & Sex Age Work Exp. Team (Industry) Team Size
W1 Worker, Female 28 1–2 years Product Team (Design) 10-15
M2 Manager, Male 25 1–2 years Risk Management Team (Renewable Energy) 20-25
W3 Worker, Female 24 1–2 years Design Team (Automotive Manufacturing) 20-25
M4 Manager, Female 29 3–5 years Design Team (Automotive Manufacturing) 10-15
M5 Manager, Male 25 3–5 years Technical Production Team (Television Production) 30-35
M6 Manager, Male 24 1–2 years Event Planning Team (Media) 5-10
W7 Worker, Female 24 0–1 year Data Production Team (Internet) 10-15
W8 Worker, Female 26 0–1 year Research Team (Technology) 5-10
M9 Manager, Female 30 1–2 years Product Team (Education) 5-10
M10 Manager, Male 28 3–5 years Cross-Functional Team (Internet) 5-10
W11 Worker, Male 25 0–1 year Research Team (Technology and Education) ¡5
W12 Worker, Female 26 0–1 year Product Team (Internet) 5-10
M13 Manager, Female 27 1–2 years Project Execution Team (Government Agencies) ¡5
W14 Worker, Male 28 3–5 years Editorial Team (Digital Media) 5-10
W15 Worker, Male 27 3–5 years Design Team (Healthcare) 5-10
M16 Manager, Female 30 3–5 years Cross-functional Team (Professional Development) 10-15
M17 Manager, Female 31 5–10 years Marketing Team (Marketing) 5-10
M18 Manager, Male 32 5–10 years Software QA Team (Foundation Model Infrastructure) 10-15
M19 Manager, Male 45 ¿10 years Product Team (Information Technology) ¿50
W20 Worker, Female 25 1–2 years Product Team (Racing Equipment) 5-10
M21 Manager, Male 39 5–10 years Communication Design Team (Design) 5-10
M22 Manager, Female 33 5–10 years UX/UI Design Team (Technology) 10-15
W23 Worker, Female 27 3–5 years UX/UI Design Team (Design) 5-10
Our sample included 13 managers (M2, M4, M5–6, M9–10, M13, M16–19, M21–22) and 10 workers (W1, W3, W7–8, W11–12, W14–15, W20, W23). Participants ranged in age from 24 to 45 years across sectors. Notably, some workers in our sample had relatively long professional experience, while some managers were comparatively early in their careers. This reflected the fact that managerial status was not strictly tied to years of employment but to the nature of participants’ current responsibilities. In several fast-paced industries represented in our sample, such as internet and media, employees may become team leads or middle managers within one to two years, especially in organizations where project-based coordination or client-facing decision-making is delegated early. By contrast, we also included managers with extensive work experience (e.g., M18, M19, M21, M22), whose teams operated within larger or more hierarchical structures; in these settings, managerial roles emerge through tenure, accumulated expertise, and formal promotion systems. Our dataset therefore reflects a wide spectrum of managerial trajectories, from early-career managers to highly seasoned leaders, as well as workers whose long-term expertise does not necessarily translate into supervisory roles.
In this context, “work experience” refers to total years in the workforce, not the number of years spent in managerial positions. This distinction helps explain why some participants with relatively short work histories were nonetheless categorized as managers, they held formal authority in their current roles, while some participants with longer histories remained individual contributors because their duties did not involve supervising others. For this reason, role categorization in our study followed participants’ self-identification based on formal position and everyday responsibilities rather than tenure alone.
To further strengthen diversity, we also recruited participants from teams of varying sizes, ranging from small groups of fewer than five members to large teams of more than fifty. Including participants from this range allowed us to capture how the imagined role of Manager Clone Agents might differ across organizational structures and accountability demands.
During the workshops, participants were asked to fully inhabit the managerial or worker roles they currently held within their organizations, drawing on their lived experiences to co-create and reflect on speculative futures involving Manager Clone Agents. This emphasis on role-grounded participation ensured that our discussions and scenarios remained contextually rich and sensitive to the diverse ways managerial authority and worker agency are enacted across contemporary workplaces.